OpenStudy (anonymous):

In the xy plane, the point (6,3) is the midpoint of the line segment with endpoints (x,5) and (9,y) What is the value of x+y?

OpenStudy (cybershadow):

well a midpoint of a line having (x1,y1) and (x2,y2) as the vertices is (x1+x2/2,y1+y2/2) So in our case midpoint is (6,3) and end points are (x,5) and (9,y) so clearly 6 = x+9 and 3 = y+5 therefore x = 6-9 and y = 3-5 x = -3 and y = -2 and Boom! Game over .-. u have ur answer sir

OpenStudy (anonymous):

thank you so much !!

OpenStudy (cybershadow):

No worries .-.